# Characterizing the Burst Error Correction Ability of Quantum Cyclic Codes

### v1

@@ 00 $[[13,1]]$ $3$ $g=(\textbf{1}^{6}\textbf{2}^{5}\textbf{3}^{3}\textbf{2}^{1}\textbf{1}^{0})$
$[[95,19]]$ $19$ $g=(\textbf{1}^{38}\textbf{3}^{19}\textbf{1}^{0})$
$1$ $[[45,8]]$ $9$ $g_{1}=(\textbf{1}^{19}\textbf{1}^{18}\textbf{1}^{16}\textbf{1}^{12}\textbf{1}^{10}\textbf{1}^{9}\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{3}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{18}\textbf{1}^{15}\textbf{1}^{12}\textbf{1}^{9}\textbf{1}^{0})$
$2$ $[[21,9]]$ $3$ $g_{1}=(\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{1}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[23,1]]$ $5$ $g=(\textbf{1}^{11}\textbf{1}^{9}\textbf{1}^{7}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{1}\textbf{1}^{0})$
$[[31,1]]$ $7$ $g=(\textbf{1}^{15}\textbf{1}^{14}\textbf{1}^{13}\textbf{1}^{9}\textbf{1}^{8}\textbf{1}^{3}\textbf{1}^{0})$
$[[35,25]]$ $2$ $g=(\textbf{1}^{5}\textbf{2}^{4}\textbf{3}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[35,17]]$ $4$ $g=(\textbf{1}^{9}\textbf{3}^{7}\textbf{3}^{6}\textbf{3}^{5}\textbf{3}^{4}\textbf{2}^{3}\textbf{2}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[35,13]]$ $5$ $g=(\textbf{1}^{11}\textbf{3}^{10}\textbf{2}^{9}\textbf{1}^{8}\textbf{2}^{7}\textbf{2}^{6}\textbf{3}^{5}\textbf{1}^{3}\textbf{2}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[35,13]]$ $5$ $g=(\textbf{1}^{11}\textbf{3}^{10}\textbf{2}^{9}\textbf{1}^{8}\textbf{2}^{7}\textbf{2}^{6}\textbf{3}^{5}\textbf{1}^{3}\textbf{2}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[35,5]]$ $7$ $g=(\textbf{1}^{15}\textbf{1}^{14}\textbf{1}^{13}\textbf{1}^{12}\textbf{1}^{10}\textbf{1}^{8}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{4}\textbf{1}^{1}\textbf{1}^{0})$
@@
$[[63,33]]$ $7$ $g=(\textbf{1}^{15}\textbf{1}^{13}\textbf{1}^{10}\textbf{2}^{9}\textbf{2}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{2}^{5}\textbf{2}^{3}\textbf{3}^{2}\textbf{3}^{1}\textbf{3}^{0})$
$[[63,45]]$ $4$ $g=(\textbf{1}^{9}\textbf{3}^{7}\textbf{2}^{5}\textbf{2}^{4}\textbf{2}^{2}\textbf{3}^{1}\textbf{1}^{0})$
@@
$[[77,47]]$ $7$ $g=(\textbf{1}^{15}\textbf{2}^{13}\textbf{3}^{12}\textbf{1}^{11}\textbf{2}^{10}\textbf{1}^{7}\textbf{1}^{6}\textbf{2}^{5}\textbf{2}^{4}\textbf{3}^{3}\textbf{1}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[91,13]]$ $19$ $g=(\textbf{1}^{39}\textbf{1}^{38}\textbf{2}^{36}\textbf{2}^{35}\textbf{3}^{34}\textbf{3}^{32}\textbf{1}^{31}\textbf{3}^{30}\textbf{2}^{29}\textbf{2}^{27}\textbf{3}^{26}\textbf{1}^{25}\textbf{2}^{24}\textbf{3}^{23}\textbf{3}^{22}\textbf{2}^{19}\textbf{3}^{18}\textbf{3}^{17}\textbf{2}^{16}\textbf{3}^{15}\textbf{2}^{14}\textbf{1}^{13}\textbf{1}^{9}\textbf{3}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{1}^{4}\textbf{2}^{3}\textbf{2}^{2}\textbf{1}^{0})$
@@
$[[91,61]]$ $7$ $g=(\textbf{1}^{15}\textbf{1}^{13}\textbf{2}^{12}\textbf{2}^{11}\textbf{1}^{10}\textbf{1}^{9}\textbf{1}^{8}\textbf{3}^{7}\textbf{3}^{5}\textbf{3}^{2}\textbf{2}^{1}\textbf{2}^{1}\textbf{1}^{0})$
$[[91,73]]$ $4$ $g=(\textbf{1}^{9}\textbf{1}^{8}\textbf{3}^{5}\textbf{3}^{4}\textbf{1}^{3}\textbf{3}^{2}\textbf{1}^{1}\textbf{1}^{0})$
@@
Let ${C}=[n,k]_{q^{2}}$ be a classical cyclic code with a parity check matrix $\mathbf{H}$ such that $\mathbf{H}\mathbf{H}^{\dagger}=0$. Denote the generator matrix of ${C}$ by $\mathbf{G}$. Let $1\leq\ell\leq r$ and let $A_{\imath}=[\imath,\imath+\ell-1]$, where $11\leq\imath\leq n-2\ell+$ 1. Let $\mathcal{L}$ be the largest integer $\ell$ such that arbitrary one of the following two items holds
@@
    Every $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is of full rank for all $11\leq\imath\leq n-2\ell+$ 1.
@@
    For each $\imath\in[1,n-2\ell+1]$, if $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is not of full rank, there is $\mathbf{G}^{\dagger}e^{T}=\mathbf{G}^{\dagger}f^{T}$ for all $(e,f)\in\boxplus_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$.
There exists an $[[n,2k-n]]$ QCC ${Q}$ which can correct any quantum burst error of length $\mathcal{L}$ or less.
@@
In order to determine the burst error correction limit of QBECCs, we need to verify whether each $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is of full rank for $11\leq\imath\leq n-2\ell+$ 1 according to Theorem 2. If some $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is not of full rank, we need to find the set $\boxplus_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$ to verify whether the corresponding burst errors are degenerate or not. Recall that we can transform $\mathcal{M}^{(\ell)}_{A_{\imath}}$ to a diagonal matrix with form in Eq. (6) according to Lemma 5. We can conduct Gaussian elimination by rows to realize the transformation. Then the set $\boxplus_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$ can be derived directly by the “$X$” part in Eq. (6).
@@
In Table I and Table II, we list the optimal and nearly optimal QCCs of length $n<100$ by using Algorithm 1 to compute their burst error correction abilities. Denote the parameters of QCCs by $Q=[[n,k;\mathcal{L}]]$, and denote by $\Delta=n-k-4\mathcal{L}$. In Table I, we list optimal or nearly optimal nondegenerate QCCs. While in Table II, we list optimal or nearly optimal degenerate QCCs which have better burst error correction abilities than any nondegenerate QCCs constructed from the CSS or the Hermitian constructions. In Table I and Table II, the bold numbers “$\textbf{1}-\textbf{3}$” in subscripts and the numbers in superscripts of the generator polynomials stand for the coefficients and the exponents, respectively. In Table II, we denote $\ell_{0}$ by the nondegenerate burst error correction ability of the $[[n,k;\mathcal{L}]]$ QCC. In this paper, we run all the algorithms in Magma software (V2.28-3). The operating system is Ubuntu 22.04 LTS and the processor is Intel i5-12490F.
@@
26:         ///// / Do Gaussian elimination to $\mathcal{M}^{(\ell)}_{A_{\imath}}$.
27:         ~$\widetilde{\mathcal{M}}^{(\ell)}_{A_{\imath}}=$ RowsGaussianElimination$(\mathcal{M}^{(\ell)}_{A_{\imath}})$;
28:         if rank(~$(\widetilde{\mathcal{M}}^{(\ell)}_{A_{\imath}})<\ell$ then
@@
TABLE II: Computer Searching for Optimal and Nearly Optimal Nondegenerate QCCs of Length $n<100$.
[…]
00 $[[25,1]]$ $6$ $5$ $g=(\textbf{1}^{12}\textbf{2}^{11}\textbf{1}^{10}\textbf{2}^{7}\textbf{3}^{6}\textbf{2}^{5}\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[29,1]]$ $7$ $6$ $g=(\textbf{1}^{14}\textbf{2}^{13}\textbf{2}^{11}\textbf{3}^{10}\textbf{1}^{9}\textbf{3}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{1}^{5}\textbf{3}^{4}\textbf{2}^{3}\textbf{2}^{1}\textbf{1}^{0})$
$[[37,1]]$ $9$ $8$ $g=(\textbf{1}^{18}\textbf{2}^{17}\textbf{1}^{16}\textbf{1}^{15}\textbf{2}^{14}\textbf{2}^{13}\textbf{3}^{12}\textbf{1}^{11}\textbf{2}^{10}\textbf{1}^{9}\textbf{2}^{8}\textbf{1}^{7}\textbf{3}^{6}\textbf{2}^{5}\textbf{2}^{4}\textbf{1}^{3}\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[41,1]]$ $10$ $9$ $g=(\textbf{1}^{20}\textbf{2}^{19}\textbf{1}^{18}\textbf{1}^{17}\textbf{3}^{16}\textbf{3}^{14}\textbf{2}^{13}\textbf{2}^{12}\textbf{2}^{11}\textbf{3}^{10}\textbf{2}^{9}\textbf{2}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{3}^{4}\textbf{1}^{3}\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
@@
$[[85,1]]$ $21$ $20$ $g=(\textbf{1}^{42}\textbf{3}^{41}\textbf{2}^{39}\textbf{1}^{38}\textbf{1}^{37}\textbf{2}^{36}\textbf{3}^{35}\textbf{3}^{34}\textbf{3}^{33}\textbf{1}^{31}\textbf{1}^{30}\textbf{3}^{29}\textbf{3}^{28}\textbf{2}^{27}\textbf{1}^{23}\textbf{1}^{22}\textbf{3}^{21}\textbf{1}^{20}\textbf{1}^{19}\textbf{2}^{15}\textbf{3}^{14}\textbf{3}^{13}\textbf{1}^{12}\textbf{1}^{11}\textbf{3}^{9}\textbf{3}^{8}\textbf{3}^{7}\textbf{2}^{6}\textbf{1}^{5}\textbf{1}^{4}\textbf{2}^{3}\textbf{3}^{1}\textbf{1}^{0})$
$[[87,3]]$ $21$ $18$ $g=(\textbf{1}^{42}\textbf{3}^{39}\textbf{3}^{33}\textbf{2}^{30}\textbf{1}^{27}\textbf{2}^{24}\textbf{3}^{21}\textbf{2}^{18}\textbf{1}^{15}\textbf{2}^{12}\textbf{3}^{9}\textbf{3}^{3}\textbf{1}^{0})$
$[[89,1]]$ $22$ $20$
@@
$1$ $[[51,2]]$ $12$ $9$ $g_{1}=(\textbf{1}^{25}\textbf{1}^{24}\textbf{1}^{16}\textbf{1}^{15}\textbf{1}^{13}\textbf{1}^{12}\textbf{1}^{10}\textbf{1}^{9}\textbf{1}^{1}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{24}\textbf{1}^{21}\textbf{1}^{18}\textbf{1}^{12}\textbf{1}^{6}\textbf{1}^{3}\textbf{1}^{0})$
$[[85,4]]$ $20$ $15$ $g_{1}=(\textbf{1}^{41}\textbf{1}^{40}\textbf{1}^{26}\textbf{1}^{25}\textbf{1}^{21}\textbf{1}^{20}\textbf{1}^{16}\textbf{1}^{15}\textbf{1}^{1}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{40}\textbf{1}^{35}\textbf{1}^{30}\textbf{1}^{20}\textbf{1}^{10}\textbf{1}^{5}\textbf{1}^{0})$
$2$ $[[35,1]]$ $8$ $7$ $g=(\textbf{1}^{17}\textbf{1}^{15}\textbf{1}^{14}\textbf{3}^{10}\textbf{3}^{8}\textbf{3}^{7}\textbf{1}^{3}\textbf{1}^{1}\textbf{1}^{0})$
$[[79,1]]$ $19$ $18$ $g=(\textbf{1}^{39}\textbf{1}^{36}\textbf{1}^{35}\textbf{1}^{31}\textbf{1}^{30}\textbf{1}^{29}\textbf{1}^{27}\textbf{1}^{26}\textbf{1}^{25}\textbf{1}^{24}\textbf{1}^{21}\textbf{1}^{20}\textbf{1}^{19}\textbf{1}^{18}\textbf{1}^{16}\textbf{1}^{14}\textbf{1}^{13}\textbf{1}^{11}\textbf{1}^{5}\textbf{1}^{4}\textbf{1}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[93,3]]$ $22$ $20$ $g=(\textbf{1}^{45}\textbf{1}^{44}\textbf{1}^{43}\textbf{1}^{41}\textbf{1}^{40}\textbf{1}^{39}\textbf{1}^{38}\textbf{1}^{37}\textbf{1}^{34}\textbf{1}^{33}\textbf{1}^{32}\textbf{1}^{29}\textbf{1}^{28}\textbf{1}^{25}\textbf{1}^{23}\textbf{1}^{20}\textbf{1}^{19}\textbf{1}^{17}\textbf{1}^{15}\textbf{1}^{13}\textbf{1}^{11}\textbf{1}^{10}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{4}\textbf{1}^{1}\textbf{1}^{0})$

### v2

@@ 0 $[[13,1]]$ 33 $g=(\textbf{1}^{6}\textbf{2}^{5}\textbf{3}^{3}\textbf{2}^{1}\textbf{1}^{0})$
$[[95,19]]$ 1919 $g=(\textbf{1}^{38}\textbf{3}^{19}\textbf{1}^{0})$
11 $[[45,8]]$ 99 $g_{1}=(\textbf{1}^{19}\textbf{1}^{18}\textbf{1}^{16}\textbf{1}^{12}\textbf{1}^{10}\textbf{1}^{9}\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{3}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{18}\textbf{1}^{15}\textbf{1}^{12}\textbf{1}^{9}\textbf{1}^{0})$
22 $[[21,9]]$ 33 $g_{1}=(\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{1}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[23,1]]$ 55 $g=(\textbf{1}^{11}\textbf{1}^{9}\textbf{1}^{7}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{1}\textbf{1}^{0})$
$[[31,1]]$ 77 $g=(\textbf{1}^{15}\textbf{1}^{14}\textbf{1}^{13}\textbf{1}^{9}\textbf{1}^{8}\textbf{1}^{3}\textbf{1}^{0})$
$[[35,25]]$ 22 $g=(\textbf{1}^{5}\textbf{2}^{4}\textbf{3}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[35,17]]$ 44 $g=(\textbf{1}^{9}\textbf{3}^{7}\textbf{3}^{6}\textbf{3}^{5}\textbf{3}^{4}\textbf{2}^{3}\textbf{2}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[35,13]]$ 55 $g=(\textbf{1}^{11}\textbf{3}^{10}\textbf{2}^{9}\textbf{1}^{8}\textbf{2}^{7}\textbf{2}^{6}\textbf{3}^{5}\textbf{1}^{3}\textbf{2}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[35,5]]$ 77 $g=(\textbf{1}^{15}\textbf{1}^{14}\textbf{1}^{13}\textbf{1}^{12}\textbf{1}^{10}\textbf{1}^{8}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{4}\textbf{1}^{1}\textbf{1}^{0})$
@@
$[[63,33]]$ 77 $g=(\textbf{1}^{15}\textbf{1}^{13}\textbf{1}^{10}\textbf{2}^{9}\textbf{2}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{2}^{5}\textbf{2}^{3}\textbf{3}^{2}\textbf{3}^{1}\textbf{3}^{0})$
$[[63,45]]$ 44 $g=(\textbf{1}^{9}\textbf{3}^{7}\textbf{2}^{5}\textbf{2}^{4}\textbf{2}^{2}\textbf{3}^{1}\textbf{1}^{0})$
@@
$[[77,47]]$ 77 $g=(\textbf{1}^{15}\textbf{2}^{13}\textbf{3}^{12}\textbf{1}^{11}\textbf{2}^{10}\textbf{1}^{7}\textbf{1}^{6}\textbf{2}^{5}\textbf{2}^{4}\textbf{3}^{3}\textbf{1}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[91,13]]$ 1919 $g=(\textbf{1}^{39}\textbf{1}^{38}\textbf{2}^{36}\textbf{2}^{35}\textbf{3}^{34}\textbf{3}^{32}\textbf{1}^{31}\textbf{3}^{30}\textbf{2}^{29}\textbf{2}^{27}\textbf{3}^{26}\textbf{1}^{25}\textbf{2}^{24}\textbf{3}^{23}\textbf{3}^{22}\textbf{2}^{19}\textbf{3}^{18}\textbf{3}^{17}\textbf{2}^{16}\textbf{3}^{15}\textbf{2}^{14}\textbf{1}^{13}\textbf{1}^{9}\textbf{3}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{1}^{4}\textbf{2}^{3}\textbf{2}^{2}\textbf{1}^{0})$
@@
$[[91,61]]$ 77 $g=(\textbf{1}^{15}\textbf{1}^{13}\textbf{2}^{12}\textbf{2}^{11}\textbf{1}^{10}\textbf{1}^{9}\textbf{1}^{8}\textbf{3}^{7}\textbf{3}^{5}\textbf{3}^{2}\textbf{2}^{1}\textbf{2}^{1}\textbf{1}^{0})$
$[[91,73]]$ 44 $g=(\textbf{1}^{9}\textbf{1}^{8}\textbf{3}^{5}\textbf{3}^{4}\textbf{1}^{3}\textbf{3}^{2}\textbf{1}^{1}\textbf{1}^{0})$
@@
Let ${C}=[n,k]_{q^{2}}$ be a classical cyclic code with a parity check matrix $\mathbf{H}$ such that $\mathbf{H}\mathbf{H}^{\dagger}=0$. Denote the generator matrix of C{C} by $\mathbf{G}$. Let $1\leq\ell\leq r$ and let $A_{\imath}=[\imath,\imath+\ell-1]$, where $1\leq\imath\leq n-2\ell+1$. Let $\mathcal{L}$ be the largest integer $\ell$ such that exactly one of the following two terms is satisfied
@@
    Every $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is of full rank for all $1\leq\imath\leq n-2\ell+1$.
@@
    For each $\imath\in[1,n-2\ell+1]$, if $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is not of full rank, then the condition $\mathbf{G}^{\dagger}e^{T}=\mathbf{G}^{\dagger}f^{T}$ holds for all $(e,f)\in\boxplus_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$.
There exists an [[n,2kn]][[n,2k-n]] QCC Q{Q} which can correct any quantum burst error of length $\mathcal{L}$ or less.
@@
In order to determine the burst error correction limit of QBECCs, we need to verify whether each $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is of full rank for $1\leq\imath\leq n-2\ell+1$ according to Theorem 2. If some $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is not of full rank, we need to find the set $\boxplus_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$ to verify whether the corresponding burst errors are degenerate or not. Recall that we can transform $\mathcal{M}^{(\ell)}_{A_{\imath}}$ to a diagonal matrix with form in Eq. (6) according to Lemma 5. We can conduct Gaussian elimination by rows to realize the transformation. Then the set $\boxplus_{\mathcal{M}^{(\ell)}_{A_{\imath}}}$ can be derived directly by the “XX” part in Eq. (6).
@@
In Table I and Table II, we list the optimal and nearly optimal QCCs of length $n<100$ by using Algorithm 1 to compute their burst error correction abilities. Denote the parameters of QCCs by $Q=[[n,k;\mathcal{L}]]$, and denote by $\Delta=n-k-4\mathcal{L}$. In Table I, we list optimal or nearly optimal nondegenerate QCCs. While in Table II, we list optimal or nearly optimal degenerate QCCs which have better burst error correction abilities than any nondegenerate QCCs constructed from the CSS or the Hermitian constructions. In Table I and Table II, the bold numbers $\textbf{1}-\textbf{3}$” in subscripts and the numbers in superscripts of the generator polynomials stand for the coefficients and the exponents, respectively. In Table II, we denote $\ell_{0}$ by the nondegenerate burst error correction ability of the $n,k;\mathcal{L}]]$ QCC. In this paper, we run all the algorithms in Magma software (V2.28-3). The operating system is Ubuntu 22.04 LTS and the processor is Intel i5-12490F.
@@
26:   //// Do Gaussian elimination to $\mathcal{M}^{(\ell)}_{A_{\imath}}$.
[…]
@@
TABLE II: Computer searching for optimal and nearly optimal degenerate QCCs of length $n<100$.
$\Delta$ $[[n,k]]$ $\mathcal{L}$ $\ell_{0}$ Generator Polynomials
0 $[[25,1]]$ 66 55 $g=(\textbf{1}^{12}\textbf{2}^{11}\textbf{1}^{10}\textbf{2}^{7}\textbf{3}^{6}\textbf{2}^{5}\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[29,1]]$ 77 66 $g=(\textbf{1}^{14}\textbf{2}^{13}\textbf{2}^{11}\textbf{3}^{10}\textbf{1}^{9}\textbf{3}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{1}^{5}\textbf{3}^{4}\textbf{2}^{3}\textbf{2}^{1}\textbf{1}^{0})$
$[[37,1]]$ 99 88 $g=(\textbf{1}^{18}\textbf{2}^{17}\textbf{1}^{16}\textbf{1}^{15}\textbf{2}^{14}\textbf{2}^{13}\textbf{3}^{12}\textbf{1}^{11}\textbf{2}^{10}\textbf{1}^{9}\textbf{2}^{8}\textbf{1}^{7}\textbf{3}^{6}\textbf{2}^{5}\textbf{2}^{4}\textbf{1}^{3}\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
$[[41,1]]$ $10$ 99 $g=(\textbf{1}^{20}\textbf{2}^{19}\textbf{1}^{18}\textbf{1}^{17}\textbf{3}^{16}\textbf{3}^{14}\textbf{2}^{13}\textbf{2}^{12}\textbf{2}^{11}\textbf{3}^{10}\textbf{2}^{9}\textbf{2}^{8}\textbf{2}^{7}\textbf{3}^{6}\textbf{3}^{4}\textbf{1}^{3}\textbf{1}^{2}\textbf{2}^{1}\textbf{1}^{0})$
@@
$[[85,1]]$ 2121 2020 $g=(\textbf{1}^{42}\textbf{3}^{41}\textbf{2}^{39}\textbf{1}^{38}\textbf{1}^{37}\textbf{2}^{36}\textbf{3}^{35}\textbf{3}^{34}\textbf{3}^{33}\textbf{1}^{31}\textbf{1}^{30}\textbf{3}^{29}\textbf{3}^{28}\textbf{2}^{27}\textbf{1}^{23}\textbf{1}^{22}\textbf{3}^{21}\textbf{1}^{20}\textbf{1}^{19}\textbf{2}^{15}\textbf{3}^{14}\textbf{3}^{13}\textbf{1}^{12}\textbf{1}^{11}\textbf{3}^{9}\textbf{3}^{8}\textbf{3}^{7}\textbf{2}^{6}\textbf{1}^{5}\textbf{1}^{4}\textbf{2}^{3}\textbf{3}^{1}\textbf{1}^{0})$
$[[87,3]]$ 2121 $18$ $g=(\textbf{1}^{42}\textbf{3}^{39}\textbf{3}^{33}\textbf{2}^{30}\textbf{1}^{27}\textbf{2}^{24}\textbf{3}^{21}\textbf{2}^{18}\textbf{1}^{15}\textbf{2}^{12}\textbf{3}^{9}\textbf{3}^{3}\textbf{1}^{0})$
$[[89,1]]$ $22$ 2020
@@
11 $[[51,2]]$ $12$ 99 $g_{1}=(\textbf{1}^{25}\textbf{1}^{24}\textbf{1}^{16}\textbf{1}^{15}\textbf{1}^{13}\textbf{1}^{12}\textbf{1}^{10}\textbf{1}^{9}\textbf{1}^{1}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{24}\textbf{1}^{21}\textbf{1}^{18}\textbf{1}^{12}\textbf{1}^{6}\textbf{1}^{3}\textbf{1}^{0})$
$[[85,4]]$ 2020 $15$ $g_{1}=(\textbf{1}^{41}\textbf{1}^{40}\textbf{1}^{26}\textbf{1}^{25}\textbf{1}^{21}\textbf{1}^{20}\textbf{1}^{16}\textbf{1}^{15}\textbf{1}^{1}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{40}\textbf{1}^{35}\textbf{1}^{30}\textbf{1}^{20}\textbf{1}^{10}\textbf{1}^{5}\textbf{1}^{0})$
22 $[[35,1]]$ 88 77 $g=(\textbf{1}^{17}\textbf{1}^{15}\textbf{1}^{14}\textbf{3}^{10}\textbf{3}^{8}\textbf{3}^{7}\textbf{1}^{3}\textbf{1}^{1}\textbf{1}^{0})$
$[[79,1]]$ 1919 $18$ $g=(\textbf{1}^{39}\textbf{1}^{36}\textbf{1}^{35}\textbf{1}^{31}\textbf{1}^{30}\textbf{1}^{29}\textbf{1}^{27}\textbf{1}^{26}\textbf{1}^{25}\textbf{1}^{24}\textbf{1}^{21}\textbf{1}^{20}\textbf{1}^{19}\textbf{1}^{18}\textbf{1}^{16}\textbf{1}^{14}\textbf{1}^{13}\textbf{1}^{11}\textbf{1}^{5}\textbf{1}^{4}\textbf{1}^{2}\textbf{1}^{1}\textbf{1}^{0})$
$[[93,3]]$ $22$ 2020 $g=(\textbf{1}^{45}\textbf{1}^{44}\textbf{1}^{43}\textbf{1}^{41}\textbf{1}^{40}\textbf{1}^{39}\textbf{1}^{38}\textbf{1}^{37}\textbf{1}^{34}\textbf{1}^{33}\textbf{1}^{32}\textbf{1}^{29}\textbf{1}^{28}\textbf{1}^{25}\textbf{1}^{23}\textbf{1}^{20}\textbf{1}^{19}\textbf{1}^{17}\textbf{1}^{15}\textbf{1}^{13}\textbf{1}^{11}\textbf{1}^{10}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{4}\textbf{1}^{1}\textbf{1}^{0})$